OpenStudy (anonymous):

Probability: exponential random variable. Profit on a computer is $350. Suppose a warranty for one year, and the time between two failures of a computer is exponential with mean 18 months. If it costs $40 to repair a failed computer, what is the expected profit per computer? (Set up to follow)

OpenStudy (anonymous):

(Sorry to repost) Let X be an exponential random variable denoting the time until failure of a computer. E[X] = 18 months, therefore λ = 1/E[X] = 1/18 Warranty repairs must be made when X < 12 P (X<12) = 1-e-λ*12 = 1 - 1/e^(2⁄3) ≈ .48658 So, I think it should be Profit = $350 - $40 (Probability of warranty), but others have said that it's just Profit = $350 - $40*2/3; where 2/3 is the expected failure. Can anyone help clarify for me?

OpenStudy (zarkon):

I like your answer better

OpenStudy (anonymous):

Thanks for the response. Hopefully my prof will too.

OpenStudy (zarkon):

obviously we have the 350 as being fixed.. The expected loss for the warranty is \[E(L)=-40\cdot P(W\leq 12)+0\cdot P(W>12)\] \[=-40(1-e^{-2/3})+0\approx19.46\] Thus the expected profit is 350-19.46=330.54

OpenStudy (anonymous):

Super. That's what I got as well. Couldn't talk myself into 2/3, just wanted to make sure I wasn't missing something. Appreciate the effort.